# GateLens: A Reasoning-Enhanced LLM Agent for Automotive Software Release Analytics

### v1

@@
Abstract.
Ensuring the reliability and effectiveness of software release decisions is critical, particularly in safety-critical domains like automotive systems. Precise analysis of release validation data, often presented in tabular form, plays a pivotal role in this process. However, traditional methods that rely on manual analysis of extensive test datasets and validation metrics are prone to delays and high costs. Large Language Models (LLMs) offer a promising alternative but face challenges in analytical reasoning, contextual understanding, handling out-of-scope queries, and processing structured test data consistently; limitations that hinder their direct application in safety-critical scenarios.
This paper introduces GateLens, an LLM-based tool for analyzing tabular data in the automotive domain. GateLens translates natural language queries into Relational Algebra (RA) expressions and then generates optimized Python code. It outperforms the baseline system on benchmarking datasets, achieving higher F1 scores and handling complex and ambiguous queries with greater robustness. Ablation studies confirm the critical role of the RA module, with performance dropping sharply when omitted.
Industrial evaluations reveal that GateLens reduces analysis time by over 80% while maintaining high accuracy and reliability. As demonstrated by presented results, GateLens achieved high performance without relying on few-shot examples, showcasing strong generalization across various query types from diverse company roles. Insights from deploying GateLens with a partner automotive company offer practical guidance for integrating AI into critical workflows such as release validation. Results show that by automating test result analysis, GateLens enables faster, more informed, and dependable release decisions, and can thus advance software scalability and reliability in automotive systems.
@@ Abstract.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; June 23–27, 2025; Trondheim, Norwaybooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwayisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
The software industry’s transition from manual to automated processes has entered a new era with the emergence of Large Language Models (LLMs) (Liu et al., 2024; Chang et al., 2024). Companies are increasingly integrating these AI agents into their workflows, seeking more cost-effective and optimized solutions for complex software engineering tasks (Leung and Murphy, 2023). However, the direct application of LLMs to the software release validation process is hindered by limitations in interpretable reasoning and understanding of technical specifications (Marques, 2024; Austin et al., 2021).
To address these challenges, we introduce GateLens, a reasoning-enhanced LLM agent to support release validation in the automotive domain. GateLens integrates structured relational analysis with domain-specific expertise, leveraging a reasoning layer built on Relational Algebra (RA) to break down complex validation tasks into systematic analytical steps. A domain-specific knowledge base—including automotive software specifications and data schemas—improves accuracy. GateLens simplifies three critical aspects of release validation:
@@
This work aims to streamline release validation by automating key analysis workflows, enabling engineers to focus on high-value analysis and discussion. By providing deeper analytical insights, the proposed approach reduces the time needed to deliver accurate validation results, empowering release managers to make informed decisions more efficiently.
Our contributions can be summarized as follows:
  • Integrating a novel approach that combines domain-specific relational modeling with an analytical reasoning mechanism based on RA, significantly enhancing the reasoning capabilities of LLMs when working with tabular data. By leveraging RA, the approach introduces analytical steps as RA expressions, enabling precise handling of domain-specific queries without any fine-tuning.
  • A systematic framework for automotive software release validation that builds upon our existing deployed LLM-based multi-agent system in production. While our current system serves core release validation needs, increasing user diversity and query complexity revealed opportunities for improvement. GateLens addresses these limitations through enhanced query processing capabilities, as demonstrated in our comparative experiments, ensuring more robust and scalable test result analysis while maintaining the reliability of the release process.
  • An optimized system architecture for time-critical operations, minimizing LLM invocations while maintaining high accuracy. By using only a single reasoning layer or one additional invocation of the LLM, the system is lighter than approaches requiring multiple invocations for planning and execution, reducing analysis time and enhancing decision support for release managers.
  • Empirical evaluation and industrial deployment demonstrate GateLens’s effectiveness in real-world automotive software release scenarios, and that it can support diverse groups of stakeholders. Performance comparisons with GPT-4o and Llama 3.1 70B models and ablation of key system modules further highlight design trade-offs and key performance contributors.
Figure 1. GateLens top-level architecture: The system processes high-level queries from the end user, generates the necessary data manipulation code using the enhanced reasoning layer with the help of RA, executes it, and outputs the result table as a decision-support resource.
2. Background and Motivation
@@
Release managers function as gatekeepers in the software deployment pipeline. Their responsibilities encompass test result analysis, cross-system impact assessment, decision-making, stakeholder coordination, and safety compliance verification. The manual workflow introduces vulnerabilities: time-intensive processing, potential errors in interpretation, decision delays, and communication barriers between technical and business teams.
@@
3. Approach and Methodology
The complexity of automotive software release validation demands a system that can bridge the gap between human-centric inquiries and precise technical analysis. GateLens addresses this challenge by utilizing LLM agents to transform natural language queries into actionable insights through systematic analysis.
At its core, GateLens must fulfill three fundamental requirements:
@@ 3.2. Core Components
3.2.1. Query Interpreter
The query interpreter agent is responsible for converting user queries into formal RA expressions, providing a precise framework for analytical reasoning. Before initiating this translation, the agent consults the knowledge base, comprising the data schema and domain-specific context. The data schema provides a detailed understanding of the dataset, including its relational modeling, detailed field descriptions, and data types. Using this information, the agent verifies whether the query is relevant and within the scope of the dataset (Manik et al., 2021). This validation step ensures that only supported and meaningful queries are processed, improving both accuracy and efficiency. Once the query is confirmed to be in scope, the agent leverages the knowledge base to interpret and decompose the query into formal RA expressions.
@@ 3.2.1. Query Interpreter
3.2.2. Coder
Figure 2. Overview of the prompt engineering flow and prompt structure within the GateLens system architecture.
@@ 3.2.2. Coder
The generated code is immediately ready for execution and incorporates robust error handling and validation mechanisms. These include data type validation to ensure correct handling of numerical and categorical fields, null value handling to maintain data integrity, validation of numerical operations, and validation of join conditions to ensure proper matching of key columns between tables. Exception handling is also included to address potential runtime errors, ensuring reliable execution.
By generating the entire code in a single pass rather than through iterative refinement, the agent significantly reduces processing overhead, system response time, and resource consumption while minimizing potential errors that could arise from multiple execution steps. This streamlined approach to code generation and execution ensures both efficiency and reliability in the analysis pipeline while maintaining the precision established by the formal RA expressions. Producing the executable code in a single step also significantly accelerates the overall analysis process, improving system responsiveness to user queries.
3.3. Data Handling
@@
  • Hallucination Prevention: By limiting LLMs’ access to actual data, the system prevents potential hallucinations or incorrect inferences
@@
  • Knowledge Integration: Data schemas and relational models serve as a comprehensive knowledge base, providing necessary structural understanding without raw data exposure
@@ 4.1. Experimental Setup
To address the research questions introduced in Section 4, we designed and conducted extensive experiments to evaluate the performance of GateLens.
@@
The experimental data comprises two distinct benchmarks. The first benchmark consists of 50 queries designed with the assistance of release engineers, quality engineers, and verification engineers. These queries are categorized into four levels of difficulty, following the same approach introduced in GoNoGo (Khoee et al., 2024), to assess GateLens’s performance across varying levels of difficulty. The four levels of query difficulty are defined as follows:
Level 1:
@@
Simple queries involving a single operation such as filtering or sorting.
Level 2:
@@
Queries combining two or three basic operations, such as multiple filtering followed by sorting.
Level 3:
@@
Queries involving more than three operations, potentially including grouping and aggregating.
Level 4:
@@
Complex queries requiring multiple advanced operations beyond basic filtering and sorting such as grouping and aggregating for statistical calculations.
@@
The second benchmark is derived from real-world user queries collected from production logs. These queries were curated from the historical logs of the first-generation system, GoNoGo, a multi-agent system designed to support software release decision-making using planning and few-shot learning, which is currently deployed in production at our partnering industrial company (Khoee et al., 2024). While GoNoGo performs effectively in many scenarios, its limitations become evident as the range of roles and users expands, leading to a significant diversification of queries. This increased diversity exposes GoNoGo’s reliance on few-shot examples, making it less capable of handling highly complex, ambiguous, or ill-defined queries that require greater flexibility and adaptability. Despite its limitations, the GoNoGo system played a critical role in data collection for GateLens by providing query logs that aided in developing and validating future iterations. Leveraging these historical logs, we filtered out very similar queries and selected 244 of the most frequently repeated unique queries. These queries were then categorized into eight functional categories based on their purposes.
In order to assess GateLens’s performance, a series of experiments are conducted using two large language models (LLMs): GPT-4o, a leading commercial model in its class, and Llama 3.1 70b, a recently introduced open-source model. GateLens is also compared with the latest version of the first-generation system, GoNoGo (Khoee et al., 2024), our recently deployed production system that has demonstrated significant value in industrial settings. While GoNoGo effectively serves the company’s release engineering needs, this comparative analysis aims to evaluate potential improvements in our new system design. It is important to note that this version of GoNoGo differs slightly from the one introduced in the original paper, as it incorporates minor updates in prompt engineering and replaces GPT-3.5 with the aforementioned recent advanced LLMs, while still maintaining its core approach based on Chain-of-Thought (CoT) (Wei et al., 2022) reasoning and Self-Consistency (Wang et al., 2022) techniques.
@@
To address the challenge of handling out of scope user queries during real-time interactions, GateLens incorporates an in-scope filtering mechanism as explained in Section 3.2. This mechanism ensures that the system only attempts to process queries that fall within its scope, thereby improving reliability and reducing errors. Performance evaluation focused on two key aspects:
  1. (1)
@@
In our experiments, the evaluation of system performance is based on the following definitions: A True Positive (TP) occurs when the system produces a result that matches the manually generated ground-truth result. A False Positive (FP) occurs when the system provides an incorrect result. A False Negative (FN) occurs when the system fails to provide any result for a query. True Negatives (TNs) are not applicable in this evaluation, as we focus on the system’s ability to handle valid queries that result in meaningful output.
@@
4.2. Performance in Addressing User Queries (RQ1)
Table 1. Performance comparison across different models and difficulty levels on the first benchmark, which consists of 50 designed queries with annotated difficulty levels.
Level # Queries GateLens with GPT-4o GateLens with Llama 3.1 70B GoNoGo with GPT-4o GoNoGo with Llama 3.1 70B
@@ 4.2. Performance in Addressing User Queries (RQ1)
Table 2. Performance comparison of GateLens and GoNoGo across different categories on the second benchmark, which consists of 244 real-world queries.
Category # Queries GateLens with GPT-4o GateLens with Llama 3.1 70B GoNoGo with GPT-4o
@@
We conducted experiments to compare the performance of GateLens across the two introduced benchmarks. The first benchmark, consisting of 50 queries categorized by difficulty levels, was used to evaluate and compare the performance of GateLens and GoNoGo. Both systems were tested using GPT-4o and Llama 3.1 70B as their underlying LLMs. The results are summarized in Table 1.
@@
The results demonstrate that GateLens with GPT-4o significantly outperforms GateLens with Llama 3.1 70B, indicating GPT-4o’s superior capability in interpreting and generating RA. Similarly, GoNoGo with GPT-4o outperforms its Llama 3.1 70B variant, with the performance gap growing as query complexity increases. Most notably, GateLens with GPT-4o achieved optimal performance on this benchmark, maintaining 100% accuracy across all difficulty levels. This superior performance can be attributed to the RA reasoning mechanism integrated into our framework. By translating queries into RA expressions, GateLens explicitly captures the logical structure of operations, enhancing the clarity and precision of the generated code. This intermediate RA conversion allows the system to focus on the relevant table operations while filtering out irrelevant elements in the query, which greatly enhances the problem-solving capabilities of the LLM agent.
@@
For the second benchmark, results in Table 2 show that GateLens with GPT-4o and GoNoGo with GPT-4o significantly outperformed GateLens with Llama 3.1 70B. This performance disparity is primarily due to the strict code generation requirements of the task, including table filtering, merging strategies, and key-value mapping operations, where GPT-4o demonstrated markedly superior capabilities compared to Llama 3.1 70B for this purpose.
@@
GateLens with GPT-4o outperformed GoNoGo with GPT-4o across most categories, particularly evident in Metadata Queries—those seeking basic table information. For example, when processing the query ”Give me the list of release candidates,” GoNoGo often struggles with proper field identification. A common failure mode in GoNoGo occurred when user queries included typographical errors or incorrect casing in field names, with the system directly using the erroneous fields without correction. GateLens addresses this limitation through its query-to-RA transformation process, which incorporates the database’s relational model, adjusts query fields to match table formats, and can handle fuzzy matching to detect and correct field names, enabling the system to resolve typographical errors and ambiguous queries effectively. This robust approach significantly improves the system’s accuracy and resilience, particularly in real-world scenarios where user queries may not always adhere to strict formatting standards.
@@
Across both datasets, GateLens with GPT-4o consistently delivered the best performance, achieving perfect accuracy across difficulty levels in the first benchmark and superior handling of diverse query categories in the second benchmark. This superior performance can be attributed to the system’s capability to extract logical conditions from natural language queries and transform them into RA expressions. This transformation not only clarifies the underlying logic of the queries but also ensures that all conditions are accurately identified and correctly applied, minimizing errors and omissions in the generated code. These results highlight how the proposed framework significantly enhances the ability of LLMs to handle complex queries while improving their robustness and reliability in real-world applications.
4.3. Robustness: Handling Out of Scope and Imprecise Queries (RQ2)

### v2

@@
Abstract
Ensuring reliable software release decisions is critical in safety-critical domains such as automotive manufacturing. Release validation relies on large tabular datasets, yet manual analysis is slow, costly, and error-prone. While Large Language Models (LLMs) offer promising automation potential, they face challenges in analytical reasoning, structured data handling, and ambiguity resolution. This paper introduces GateLens, an LLM-based system for analyzing tabular data in the automotive domain. GateLens translates natural language queries into Relational Algebra (RA) expressions and generates optimized Python code. Unlike traditional multi-agent or planning-based systems that can be slow, opaque, and costly to maintain, GateLens emphasizes speed, transparency, and reliability. Experimental results show that GateLens outperforms the existing Chain-of-Thought (CoT) + Self-Consistency (SC) based system on real-world datasets, particularly in handling complex and ambiguous queries. Ablation studies confirm the essential role of the RA layer. Industrial deployment shows over 80% reduction in analysis time while maintaining high accuracy across test result interpretation, impact assessment, and release candidate evaluation. GateLens operates effectively in zero-shot settings without requiring few-shot examples or agent orchestration. This work advances deployable LLM system design by identifying key architectural features—intermediate formal representations, execution efficiency, and low configuration overhead—crucial for safety-critical industrial applications.
Index Terms:
@@ Abstract
I Introduction
@@
The software industry’s transition from manual to automated processes has entered a new era with the emergence of Large Language Models (LLMs) (Liu et al., 2024; Chang et al., 2024). Companies are increasingly integrating these AI agents into their workflows, seeking more cost-effective and optimized solutions for complex software engineering tasks (Leung and Murphy, 2023). However, direct application of LLMs to software release validation is hindered by limitations in interpretable reasoning and understanding of technical specifications (Marques, 2024; Austin et al., 2021).
To address these challenges, we introduce GateLens, a reasoning-enhanced LLM agent (Miehling et al., 2025) to support release validation in the automotive domain. GateLens integrates structured relational analysis with domain-specific expertise, leveraging a reasoning layer built on Relational Algebra (RA) to break down complex validation tasks into systematic analytical steps. GateLens simplifies three critical aspects of release validation:
@@
This work aims to streamline release validation by automating key analysis workflows, enabling engineers to focus on high-value analysis and discussion. By providing deeper analytical insights, the proposed approach reduces the time needed to deliver accurate validation results, empowering release managers to make informed decisions more efficiently. Our contributions can be summarized as follows:
  • We design an architecture optimized for time- and safety-critical environments, minimizing LLM invocations while preserving reasoning depth. GateLens operates in a zero-shot setting, avoiding the need for few-shot examples or multi-agent coordination, which improves generalizability, execution speed, reduces maintenance overhead, and enhances transparency.
  • We introduce a scalable and maintainable framework for automotive software release validation, developed in response to observed limitations in traditional planning-based multi-agent system. GateLens handles diverse user queries with higher robustness and clarity, supporting effective decision-making across a wider range of release engineering tasks.
  • We conduct a comprehensive empirical evaluation, including comparisons with a CoT+SC system, ablation of the RA module, and performance across multiple LLMs (GPT-4o and Llama 3.1 70B). These experiments demonstrate GateLens’s performance advantages in complex and ambiguous queries.
  • We report on real-world industrial deployment in a partner automotive company, where GateLens reduces analysis time by over 80% and demonstrates strong generalization across user roles, highlighting its practical value and deployment-readiness in safety-critical release validation workflows.
Figure 1: GateLens top-level architecture: The system processes high-level queries from the end user, generates the necessary data manipulation code using the enhanced reasoning layer with the help of RA, executes it, and outputs the result table as a decision-support resource.
II Background and Motivation
@@
Release managers function as gatekeepers in the software deployment pipeline. They handle test result analysis, cross-system impact assessment, decision-making, stakeholder coordination, and safety compliance verification. The manual workflow introduces vulnerabilities: time-intensive processing, potential errors in interpretation, decision delays, and communication barriers between technical and business teams.
@@
III Approach and Methodology
The complexity of automotive software release validation demands a system that can bridge the gap between human-centric inquiries and precise technical analysis. GateLens addresses this challenge by utilizing LLM agents to transform natural language queries into actionable insights through systematic analysis. At its core, GateLens must fulfill three fundamental requirements:
@@ III-B Core Components
III-B1 Query Interpreter
The query interpreter agent is responsible for converting user queries into formal RA expressions, providing a precise framework for analytical reasoning. Before initiating this translation, the agent consults the knowledge base, comprising the data schema and domain-specific context. The data schema provides a detailed understanding of the dataset, including its relational modeling, detailed field descriptions, and data types. Using this information, the agent verifies whether the query is relevant and within the scope of the dataset (Manik et al., 2021). This validation step ensures that only supported and meaningful queries are processed, improving both accuracy and efficiency. Once the query is confirmed to be in scope, the agent leverages the data schemas to interpret and decompose the query into formal RA expressions.
@@ III-B1 Query Interpreter
III-B2 Coder
@@
To ensure the generated code meets quality standards, we designed prompts that specifically instruct the LLM to include essential validation mechanisms. The prompts explicitly require data type validation instructions for numerical and categorical field handling, null value checking procedures to maintain data integrity, validation of numerical operations, and validation of join conditions to ensure proper matching of key columns between tables.
By generating the entire code in a single pass rather than through iterative refinement, the agent significantly reduces processing overhead, system response time, and resource consumption while minimizing potential errors that could arise from multiple execution steps. This streamlined yet precise approach ensures both efficiency and reliability in the analysis pipeline, fully aligned with the formal rigor of RA expressions and improving overall system responsiveness to user queries.
Figure 2: Overview of the prompt engineering flow and prompt structure within the GateLens system architecture.
III-C Data Handling
@@
  • Knowledge Integration: Data schemas and relational models serve as an essential knowledge base, providing necessary structural understanding without raw data exposure
@@ IV-A Experimental Setup
To address the research questions introduced in Section IV, we designed and conducted extensive experiments to evaluate the performance of GateLens.
@@
The experimental data comprises two distinct benchmarks. The first benchmark consists of 50 queries designed with the assistance of release engineers, quality engineers, and verification engineers. To assess GateLens’s performance across a spectrum of query complexities, these queries are categorized into four difficulty levels. The four levels of query difficulty are defined as follows:
@@
Level 1 Simple queries involving a single operation such as filtering or sorting.
@@
Level 2 Queries combining two or three basic operations, such as multiple filtering followed by sorting.
@@
Level 3 Queries involving more than three operations, potentially including grouping and aggregating.
@@
Level 4 Complex queries requiring multiple advanced operations beyond basic filtering and sorting, such as grouping and aggregating for statistical calculations.
@@
The second benchmark is derived from real-world user queries collected from production logs at our partner company. These queries were sourced from the historical logs of an agentic system that employed a well-established tabular data reasoning approach combining CoT (Wei et al., 2022) prompting with Self-Consistency (SC) (Wang et al., 2022). This system was used in production to support software release analytics, and the collected queries reflect a wide range of user roles, query types, and domain-specific requirements. While this system performs effectively in many scenarios, its limitations become apparent as the range of roles and users expands, leading to a significant diversification of queries. This broader query diversity exposes the system’s reliance on few-shot examples, making it less capable of handling highly complex, ambiguous, or ill-defined queries that require greater flexibility and adaptability. Nevertheless, this preliminary system played a critical role in data collection for GateLens by providing query logs used to develop and validate our approach. From these logs, we filtered out near-duplicates and selected 244 frequently repeated unique queries, which we then organized into eight functional categories based on their purposes.
In order to assess GateLens’s performance, we run experiments with two large language models (LLMs): GPT-4o, a leading commercial model, and Llama 3.1 70b, a recently released open-source model. We also benchmark GateLens against the CoT+SC agentic system currently used to support the company’s release decisions. Our comparative analysis is designed to quantify the improvements introduced by GateLens’s novel architecture.
@@
To address the challenge of handling out-of-scope user queries during real-time interactions, GateLens incorporates an in-scope filtering mechanism as explained in Section III-B. This mechanism ensures that the system only attempts to process queries that fall within its scope, thereby improving reliability and reducing errors. Performance evaluation focused on two key aspects:
  1. 1.
@@
In our experiments, the evaluation of system performance is based on the following definitions: A True Positive (TP) occurs when the system produces a result that matches the manually generated ground-truth result, specifically when the final output of the executed code matches the expected ground-truth output. A False Positive (FP) occurs when the system provides an incorrect result, meaning the executed code produces output that differs from the ground-truth. A False Negative (FN) occurs when the system fails to provide any result for a query. True Negatives (TNs) are not applicable since we focus on valid queries producing meaningful output.
@@
IV-B Performance in Addressing User Queries (RQ1)
TABLE I: Performance comparison across different models and difficulty levels on the first benchmark, which consists of 50 designed queries with annotated difficulty levels.
Level # Queries GateLens with GPT-4o GateLens with Llama 3.1 70B CoT+SC with GPT-4o CoT+SC with Llama 3.1 70B
@@ IV-B Performance in Addressing User Queries (RQ1)
TABLE II: Performance comparison of GateLens and the CoT+SC system across different categories on the second benchmark, which consists of 244 real-world queries.
Category # Queries GateLens with GPT-4o GateLens with Llama 3.1 70B CoT+SC with GPT-4o
@@
We conducted experiments to compare the performance of GateLens across the two introduced benchmarks. The first benchmark, consisting of 50 queries categorized by difficulty levels, was used to evaluate and compare the performance of GateLens and CoT+SC. Both systems were tested using GPT-4o and Llama 3.1 70B as their underlying LLMs. The results are summarized in Table I.
@@
The results demonstrate that GateLens with GPT-4o significantly outperforms GateLens with Llama 3.1 70B, indicating GPT-4o’s superior capability for interpreting and generating RA. Similarly, CoT+SC with GPT-4o outperforms its Llama 3.1 70B variant, with the performance gap growing as query complexity increases. CoT+SC performance declines with query complexity. This underscores the importance of the RA reasoning mechanism in GateLens, which enables effective handling of complex, unstructured queries by decomposing them into logical, structured expressions. Most notably, GateLens with GPT-4o achieved optimal performance on this benchmark, maintaining 100% accuracy across all difficulty levels. This stems from integrating RA reasoning into our framework. By translating queries into RA expressions, GateLens explicitly captures the logical structure of operations, enhancing both the clarity and precision of the generated code. The intermediate RA conversion allows the system to focus on the relevant table operations while filtering out irrelevant elements in the query, greatly enhancing the problem-solving capabilities of the LLM agent.
@@
For the second benchmark, results in Table II show that GateLens (GPT-4o) and CoT+SC (GPT-4o) significantly outperformed GateLens with Llama 3.1 70B. This performance disparity is primarily due to the strict code generation requirements of the task, including table filtering, merging strategies, and key-value mapping operations, where GPT-4o demonstrated markedly superior capabilities.
@@
GateLens with GPT-4o outperformed CoT+SC (GPT-4o) across most categories, particularly evident in Metadata Queries (those seeking basic table information). For example, when processing the query ”Give me the list of release candidates,” the CoT+SC system often fails to identify the correct field. A common failure mode in CoT+SC occurred when user queries included typographical errors or incorrect casing in field names, with the system directly using the erroneous fields without correction. GateLens addresses this limitation through its query-to-RA transformation process, which incorporates the database’s relational model, adjusts query fields to match table formats, and can handle fuzzy matching to detect and correct field names, enabling the system to resolve typographical errors and ambiguous queries effectively. This approach improves accuracy and resilience, particularly in real-world scenarios where user queries may not always adhere to strict formatting standards.
@@
Across both datasets, GateLens with GPT-4o consistently delivered the best performance, achieving perfect accuracy across difficulty levels in the first benchmark and superior handling of diverse query categories in the second benchmark. This superior performance stems from the system’s ability to extract logical conditions from natural language queries and transform them into RA expressions, clarifying query logic, ensuring accurate, complete condition handling and minimizing errors and omissions in the generated code. These results demonstrate the framework’s enhancement of LLM performance on complex queries and improved real-world reliability
IV-C Robustness: Handling Out of Scope and Imprecise Queries (RQ2)